OpenStudy (anonymous):

What is the directrix of the parabola with the equation y+3=1/10(x+2)squared?

OpenStudy (imstuck):

I am getting the line y = -11/2. But don't quote me on it.

OpenStudy (imstuck):

Can you test the answer to see it's incorrect? Or do you have a choice of answers?

OpenStudy (anonymous):

Possible answers are x=-3, y=-0.5, y=-5.5, and x=-2

OpenStudy (imstuck):

Ok, well -11/2 is -5.5 So it's y = -5.5

OpenStudy (imstuck):

Do you want to know how I got it?

OpenStudy (anonymous):

Than you very much. Yes please

OpenStudy (imstuck):

Ok, the equation that you gave us can be rewritten to get rid of the 1/10 fraction. Do that by multiplying the 10 by both sides to get this:\[10(y+3)=(x+2)^{2}\]On the right side the 10 and the 1/10 cancel each other out. I didn't show that step; it's understood (?)

OpenStudy (anonymous):

Thank you

OpenStudy (imstuck):

Now distribute the 10 into the parenthesis to get 10y + 30. Now the equation for finding the directrix and the focus is x^2 = 4py, where 10, the coefficient, is y. So 4p=10 and solve for p. P gives you 10/4=5/2. But since the center is at (-2,-3) and this is an x^2 parabola, the focus lies on the line of symmetry and the directrix is a y = line the same distance from the vertex as is the focus. The focus is 5/2 above the center of (-2,-3) and the directrix is 5/2 below the center of x=-2. It looks like this:

OpenStudy (imstuck):